A proton experiences an eastward force of 30 × N while traveling in a magnetic field of 5 × T north. 

  1. Find the magnitude of the proton’s velocity
  2. State the direction of the proton’s velocity 
  3. Draw the proton and indicate the directions of all three vectors using the appropriate notation as shown in the table above
1
Expert's answer
2021-05-04T07:03:21-0400

To be given in question

Magnetic force(F)=30i


Magnetic field (B)=5j

To be asked in question

Part(a) magnitude of velocity

Part(b)=velocity direction


Part (c)=diagram

Part (a)

We know that

"F=q(v\times B)"

"F=qvBsin\theta\hat{n}"

"\hat{n} direction is (v\times B) direction"

Magnitude velocity

"v=\frac{F}{qB}"

Put value

"v=\frac{30}{1.6\times10^{-19}\times5}"

"v=3.75\times 10^{19}m/sec"


Part (b) state direction of proton velocity

"F=q(v\times B)"

"F=qvB\hat{n}"

"\hat{n} direction is"

"(v\times B)" direction mean F direction (i) direction


"(v\times B)" direction (i) direction possible then velocity direction (-k)

Then velocity direction

"v=v(-k)"

Nagetive z-axis direction of velocity direction

Part(c)

Vector form daigram




According to diagram velocity direction

"(-\hat{k})"